OpenStudy (anonymous):

Trig substitution question: Integrate sqrt(4^2 -x^2) from 0 to 2. My work: x = 4sinθ x^2 = 4^2 *sin^2 θ sinθ = x/4 sqrt(4^2(1-sin^2 θ))) = sqrt(4^2 *cos^2 θ) =4cosθ integral 4cosθ = 4sinθ sub for sinθ.... x from 0 to 2. Unfortunately, this isn't right. How should I have done the problem?

OpenStudy (anonymous):

@Hero @robtobey @campbell_st

OpenStudy (anonymous):

\[\int_0^2\sqrt{4^2-x^2}\;dx\\ \text{letting }x=4\sin\theta \iff \theta=\arcsin\frac{x}{4},\\ dx=4\cos\theta\;d\theta,\\ \text{making your integral}\\ \int_0^2\sqrt{4^2-\left(4\sin\theta\right)^2}\;\left(4\cos\theta\;d\theta\right)\] Since you've changes the limit of integration, you have to change the limits, too. The interval [0,2] corresponds with x. Now, you're integrating with respect to theta. The new limits will be \[\text{Upper: } \theta=\arcsin\frac{2}{4}=\arcsin\frac{1}{2}=\frac{\pi}{6}\\ \text{Lower: }\theta = \arcsin\frac{0}{4}=0\] With the new limits, the integral is now \[\large \int_0^{\frac{\pi}{6}}\sqrt{4^2-\left(4\sin\theta\right)^2}\;\left(4\cos\theta\;d\theta\right)\\ \large 16\int_0^{\frac{\pi}{6}}\sqrt{1-\sin^2\theta}\;\cos\theta\;d\theta\]